A flower pot falls from a window sill 25.0m above the sidewalk.
a.How fast is the flowerpot moving when it strikes the ground?
b.How much time does a passerby on the sidewalk below have to move out of the way before the flower pot hits the ground?

Answer by Boreal(15235)

s=(1/2)at^2=(1/2)9.8 m/sec^2*t^2=25
50=9.8m/sec^2*t^2
t^2=5.10 sec^2't=2.26 sec ANSWER b
v=gt=9.8 m/sec*2.28 sec=22.344 m/sec ANSWER a